Find the dimension and the maximum area of a rectangle if its perimeter is 36m.
Answer by rothauserc(4718)

let l be length and w be width of the rectangle, then
2l + 2w = 36
divide both sides of = by 2
l + w = 18
the area of the rectangle will be
A = l * w
let l = 9 and w = 9 the area is
note that all squares are rectangles, the maximum area is
A = 9 * 9 = 81 square meters (m^2)